# Cognitive Hierarchy in Day-to-day Network Flow Dynamics

### v1

@@ 2.4.2 Practical relevance
Our modeling framework may be used to characterize the strategic thinking behaviors of major decision-making entities that each control a significant portion of traffic flow. In this context, the flow distribution vector of one class can be viewed as a decision of one large entity. This approach is particularly relevant in modern urban environments where traffic patterns are increasingly influenced by sophisticated navigation systems and centralized management centers. These systems operate at different levels of sophistication in predicting and responding to traffic conditions. At the most basic level ($k=0$), we have navigation systems or drivers simply reacting to current traffic conditions without strategic planning. The next level ($k=1$) includes more advanced navigation platforms like Google Maps and Waze, which not only provide route recommendations based on current conditions but also engage in predictive behavior (Lau, 2020). By anticipating the actions of other road users and potential congestion points, they aim to recommend shorter routes for their users, effectively playing a strategic game against other decision-makers in the network. At higher level of sophistication ($k=2$ or above) are advanced traffic management centers and mobility service providers. These entities attempt to optimize overall traffic flow by considering multiple levels of interaction between different road user groups. Their decisions may require anticipating the responses of various traveler segments and adjusting traffic control measures accordingly.
@@ 3 CH-NTP Dynamic
The previous section describes the idea using a general day-to-day operator $H[\cdot]$. This section specifies $H[\cdot]$ as the particular NTP dynamic that admits DUE as a fixed point and analyzes its mathematical properties. The NTP dynamic’s behavioral explanation is identical to a link-based day-to-day model proposed in He et al. (2010): travelers seek to minimize their travel costs while exerting some efforts (or incurring costs) to deviate from incumbent routes. This explanation was also discovered by Tsakas and Voorneveld (2009) in the game theory literature in which the NTP dynamic was called the target projection dynamic. In addition, the NTP dynamic has a unified closed-form formula with the projected dynamical system when route flows only evolve in the interior of the feasible route flow set (Nagurney and Zhang, 1997; Guo et al., 2015; Xiao et al., 2016).
@@
where $P_{\Omega_{\eta}}{\left[\boldsymbol{z}\right]}$ is the projection operator that solves the optimization $\operatorname*{arg\,min}_{\boldsymbol{h}\in\Omega_{\eta}}\left\lVert{\boldsymbol{h}-\boldsymbol{z}}\right\rVert^{2}$. The $\zeta$ in (1) is now replaced by $\gamma(>0)$, which captures the costs incurred by deviations from the incumbent routes. With a larger $\gamma$, travelers have smaller inertia and are prone to switching.
Based on the modeling framework in Section 2.3, the predicted flow pattern of each class $k\in\{0,1,2\}$ is given by:
@@
where the predicted parameter $\hat{\zeta}$ in (4) and (6) is replaced by $\hat{\gamma}>0$. Eqs. (10)-(12) indicate that for each $k$, $\boldsymbol{\pi}^{k,(t+1)}$ is a function of $\tilde{\boldsymbol{x}}^{(t)}$.
@@
(Facchinei and Pang, 2007). Let $C$ be a nonempty and closed convex subset of $\mathcal{R}^{n}$. Then for any $\boldsymbol{z}\in\mathcal{R}^{n}$ and $\boldsymbol{w}\in C$,
@@ 3.1 Mixed prediction-based equilibria
Fixed points of the dynamical system (9)-(12) are termed as mixed prediction-based equilibria (MPE) and they are characterized in the following proposition.
@@
Continuity of the RHS of the dynamical system (9)-(12) is guaranteed by the continuities of the projection operator (Penot, 2005) and the route cost function. Hence, according to Brouwer’s Fixed Point Theorem, there exist at least one fixed point to the dynamical system.
@@
We next define a $|K|$-class DUE in Definition 1 (see also in Nagurney, 2000, and Zhou et al., 2020) and show that it is one of the MPE in Proposition 2.
@@
(Nagurney, 2000; Zhou et al., 2020) A route flow pattern $\boldsymbol{x}^{\ast}\equiv(\boldsymbol{x}^{k,\ast},\boldsymbol{x}^{k,\ast}\in\Omega_{p^{k}},k\in K)^{T}$ is said to be a $|K|$-class DUE if the following VI holds:
@@
A $|K|(=3)$-class DUE in Definition 1 is an MPE of the dynamical system (9)-(12), but not vice versa.
@@
We prove the sufficiency, and the necessity can be easily negated by a counter-example, as demonstrated in the numerical experiments (see Section 5.1).
At the DUE, first note from (10) that $\boldsymbol{\pi}^{0,\ast}$ equals the aggregate flow pattern, $\tilde{\boldsymbol{x}}^{\ast}$. It is well-known that $\tilde{\boldsymbol{x}}^{\ast}\in\Omega_{1}$ of the classical single-class DUE satisfies the following VI:
@@
Multiplying (17) by $-\hat{\gamma}$ and adding $(\tilde{\boldsymbol{x}}^{\ast})^{T}(\boldsymbol{x}-\tilde{\boldsymbol{x}}^{\ast})$ to both sides yields:
@@
Hence, according to (11), at the DUE, the predicted flow pattern of class-1 player, $\boldsymbol{\pi}^{1,\ast}=\tilde{\boldsymbol{x}}^{\ast}$.
@@
Applying (3.1) to the definition of the 2-step travelers’ predicted cost in (12), we obtain that the predicted flow pattern of class-2 player, $\boldsymbol{\pi}^{2,\ast}=\tilde{\boldsymbol{x}}^{\ast}$.
Summarizing the above, the predicted flow patterns of all the three classes at DUE are exactly $\tilde{\boldsymbol{x}}^{\ast}$, and thus (14) becomes (16). ∎
@@ 3.2 Local stability
The MPE’s non-uniqueness makes global stability analysis difficult, if not impossible, because we cannot construct a global convex Lyapunov function that decreases as time passes. We, therefore, turn to analyze its local stability. The definitions and theorems on local stability used in this paper are relegated to the Appendix A. In this subsection, we first give the Jacobian of the CH-NTP dynamic at the MPE and then analyze the stability near the DUE.
@@ 3.2.1 Jacobian matrix
With Lemma 2, the Jacobian matrix for the CH-NTP dynamic can be derived as follows and used to examine the local stability at any MPE by checking its eigenvalues.
@@
where $D_{k}\equiv D\left[\boldsymbol{\pi}^{k,(t+1)}\right]$ representing the Jacobian of the route cost functions evaluated at $\boldsymbol{\pi}^{k,(t+1)}$; $Q^{k}\equiv Q\left[\boldsymbol{x}^{k,(t)}-\gamma\boldsymbol{c}\left(\boldsymbol{\pi}^{k,(t+1)}\right)\right]$; $I$ the identity matrix of size $\sum_{w}{|R_{w}|}$; and $\hat{Q}_{1}^{0}\equiv Q\left[q_{1}^{0}\tilde{\boldsymbol{x}}^{(t)}-\hat{\gamma}\boldsymbol{c}\left(\boldsymbol{\pi}^{0,(t+1)}\right)\right]$, $\hat{Q}_{2}^{0}\equiv Q\left[q_{2}^{0}\tilde{\boldsymbol{x}}^{(t)}-\hat{\gamma}\boldsymbol{c}\left(\boldsymbol{\pi}^{0,(t+1)}\right)\right]$, $\hat{Q}_{2}^{1}\equiv Q\left[q_{2}^{1}\tilde{\boldsymbol{x}}^{(t)}-\hat{\gamma}\boldsymbol{c}\left(\boldsymbol{\pi}^{1,(t+1)}\right)\right]$.
@@ 3.2.2 Stability around the DUE
We are particularly interested in how the travelers’ prediction behaviors would affect the stability near the DUE. Numerically, it is always tractable to calculate $JP$’s eigenvalues at the DUE and check its local stability. However, general analytical insights on the parametric space may rely on certain assumptions. Recognizing this, we assume that as the day-to-day process evolves, the projection always yields a positive flow on each route. Under this assumption, the NTP dynamic reduces to a closed-form expression (Sandholm, 2010; Xiao et al., 2016). The $\alpha$ and $\gamma$ (and $\hat{\alpha}$ and $\hat{\gamma}$) can be simply combined (specifically, multiplied) to represent the sensitivity to the cost difference between two routes; see Appendix A in Xiao et al. (2016). We can therefore fix $\alpha=\hat{\alpha}=1$ and vary $\gamma$ and $\hat{\gamma}$ while the behavioral explanation is not compromised.
We divide the discussion into two parts: (i) when higher-step travelers can exactly predict lower-step travelers’ switching tendency, i.e., $\hat{\gamma}=\gamma$; and (ii) when the prediction is inaccurate, i.e., $\hat{\gamma}\neq\gamma$.
The following assumption is made on the route cost function when analyzing the stability around the DUE.
@@ 3.2.2.1   Perfect prediction $\hat{\gamma}=\gamma$
For a dynamical system (9)-(12) featuring any $|K|\in\{1,2,3\}$, if the projection operator can always generate positive flow on each route (hence $\hat{\alpha}$ and $\alpha$ can be set to 1) and $\hat{\gamma}=\gamma$, its DUE under Assumption 1 is locally stable if all the moduli of the eigenvalues of matrix $\bar{Q}(I-\gamma D_{\ast})$ are less than 1, where $D_{\ast}$ is the Jacobian of the route cost function evaluated at the aggregate DUE, $\tilde{\boldsymbol{x}}^{\ast}$, and $\bar{Q}$ is defined in Lemma 2.
@@
The difference between models with different $|K|$ is that (as we shall see in the numerical experiments) the instability condition of $|K|=1$ leads to a permanent oscillating pattern near the DUE (i.e., crossing the DUE infinitely many times) while for $|K|\geq 2$ it is more likely that the system will evolve to and stick at an MPE that is not DUE.
@@
See Appendix E. ∎
@@
Although revealed by a simple two-route network, the above finding still holds in general networks, as shown in the following proposition.
@@
Consider a dynamical system (9)-(11) with $|K|=2$ where the projection can always generate positive flow on each route (hence $\hat{\alpha}$ and $\alpha$ can be set to 1). Denote $\beta_{i}$ as $\bar{Q}D_{\ast}$’s $i$-th eigenvalue and $\beta_{i}\geq 0,\forall i$. The DUE is locally stable if and only if $\left|{\gamma\hat{\gamma}\beta_{i}^{2}-2\gamma\beta_{i}+1}\right|<1,\forall i$.

### v2

@@ 2.4.3 Practical relevance
Our modeling framework may also be used to characterize the strategic thinking behaviors of major decision-making entities that each control a portion of traffic flow. In this context, the flow distribution vector of one class can be viewed as a decision of one large entity. This approach has particular relevance in modern transportation networks where route choices are increasingly determined by sophisticated navigation systems and centralized management centers operating at different levels of strategic sophistication:
Basic reactive systems ($k=0$): Simple navigation apps and individual drivers react to current traffic conditions without strategic anticipation.
Predictive navigation platforms ($k=1$): Advanced navigation services like Google Maps provide route recommendations based not only on current traffic conditions but also on predictive models of future traffic states (Lau 2020). Since recommending the current faster routes to too many users may create new congestion on these routes (Sun et al. 2020), navigation services might adopt a strategic approach. They anticipate the actions of other road users and potential congestion points. Based on these predictions, they recommend shorter routes for their users. Anecdotal evidence from online forums suggests the navigation app Waze may already implement such predictive traffic distribution algorithms to distribute traffic and reduce overall network congestion (Waze-user 2024, 2025). While such observations are not conclusive, they suggest that principles similar to strategic prediction may already operate in real-world navigation systems. Similarly, Gaode Maps (from China) has reported implementing what they refer to as “traffic load-balancing” when planning routes, which distributes traffic across multiple paths to prevent congestion on any single route (Gaode-Map 2020).
Strategic traffic management ($k=2$ or above): Advanced traffic management centers and mobility service providers may attempt to optimize overall traffic flow by considering multiple levels of interaction between different traveler groups. Their decisions may require anticipating the responses of these groups and adjusting traffic control measures accordingly.
@@ 3 CH-NTP Dynamic
The previous section describes the idea using a general day-to-day operator $H[\cdot]$. This section specifies $H[\cdot]$ as the particular NTP dynamic that admits DUE as a fixed point. The NTP dynamic’s behavioral explanation is identical to a link-based day-to-day model proposed in He, Guo, and Liu (2010): travelers seek to minimize their travel costs while exerting some efforts (or incurring costs) to deviate from incumbent routes. In addition, the NTP dynamic has a unified closed-form formula with the projected dynamical system when route flows only evolve in the interior of the feasible route flow set (Nagurney and Zhang 1997, Guo et al. 2015, Xiao, Yang, and Ye 2016).
@@
where $P_{\Omega_{\eta}}{\left[\boldsymbol{z}\right]}$ is the projection operator that solves the optimization $\operatorname*{arg\,min}_{\boldsymbol{h}\in\Omega_{\eta}}\left\lVert{\boldsymbol{h}-\boldsymbol{z}}\right\rVert^{2}$. The $\zeta$ in Eq. (1) is now replaced by $\gamma(>0)$, which captures the costs incurred by deviations from the incumbent routes. With a larger $\gamma$, travelers have smaller inertia and are prone to switching.
As per Section 2.3, the predicted flow pattern of each class $k\in\{0,1,2\}$ is given by:
@@
where the predicted parameter $\hat{\zeta}$ in Eqs. (4) and (6) is replaced by $\hat{\gamma}(>0)$. Eqs. (10)-(12) indicate that for each $k$, $\boldsymbol{\pi}^{k,(t+1)}$ is a function of $\tilde{\boldsymbol{x}}^{(t)}$.
@@
(Facchinei and Pang 2007). Let $C$ be a nonempty and closed convex subset of $\mathcal{R}^{n}$. Then for any $\boldsymbol{z}\in\mathcal{R}^{n}$ and $\boldsymbol{w}\in C$,
@@ 3.1 Mixed prediction-based equilibria
Fixed points of the dynamical system in Eqs. (9)-(12) are termed as mixed prediction-based equilibria (MPE) and they are characterized in the following proposition.
@@
Continuity of the RHS of the dynamical system in Eqs. (9)-(12) is guaranteed by the continuity of both the projection operator (Penot 2005) and the route cost function. According to Brouwer’s Fixed Point Theorem, at least one fixed point exists for the dynamical system.
@@
[…]
These inequalities take the form of VIs. We can combine them because sets $\Omega_{p^{k}}(k\in K)$ are mutually disjoint (Kinderlehrer and Stampacchia 2000). ∎
We next define a $|K|$-class DUE in Definition 1 (see also in Nagurney 2000, and Zhou, Xu, and Meng 2020) and show that it is one of the MPE in Proposition 2.
@@
(Nagurney 2000, Zhou, Xu, and Meng 2020) A route flow pattern $\boldsymbol{x}^{\ast}\equiv(\boldsymbol{x}^{k,\ast},\boldsymbol{x}^{k,\ast}\in\Omega_{p^{k}},k\in K)^{T}$ is said to be a $|K|$-class DUE if the following VI holds:
@@
A $|K|(=3)$-class DUE in Definition 1 is an MPE of the dynamical system in Eqs. (9)-(12), but not vice versa.
@@
We prove the sufficiency, and the necessity can be easily negated by a counter-example, as will be demonstrated in Section 3.3.2).
At the DUE, Eq. (10) indicates that $\boldsymbol{\pi}^{0,\ast}$ equals the aggregate flow pattern, $\tilde{\boldsymbol{x}}^{\ast}$. It is well-known that the classical single-class DUE $\tilde{\boldsymbol{x}}^{\ast}\in\Omega_{1}$ satisfies the following VI:
@@
We multiply Eq. (18) by $-\hat{\gamma}$. We then add $(\tilde{\boldsymbol{x}}^{\ast})^{T}(\boldsymbol{x}-\tilde{\boldsymbol{x}}^{\ast})$ to both sides and rearrange the terms. This yields:
@@
Hence, according to Eq. (11), at the DUE, the predicted flow pattern of class-1 player, $\boldsymbol{\pi}^{1,\ast}=\tilde{\boldsymbol{x}}^{\ast}$.
@@
In Eq. (3.1), the first equality applies the definition of projection. The second equality occurs because dividing the objective by a constant $\eta$ does not affect the optimal solution. The third equality holds because the optimization problem commutes with scaling of the demand ($\eta$) due to the linear structure of $\Omega_{\eta}$. The fourth equality follows from Eq. (20).
Applying Eq. (3.1) to the definition of the 2-step travelers’ predicted cost in Eq. (12), we obtain that the predicted flow pattern of class-2 player, $\boldsymbol{\pi}^{2,\ast}=\tilde{\boldsymbol{x}}^{\ast}$.
Summarizing the above, the predicted flow patterns of all the three classes at DUE are exactly $\tilde{\boldsymbol{x}}^{\ast}$, and thus Eq. (14) becomes Eq. (17). ∎
@@ 3.2 Local stability
The MPE’s non-uniqueness makes global stability analysis difficult, if not impossible, because we cannot construct a global convex Lyapunov function that decreases as time passes. We therefore focus on analyzing its local stability. The definitions and theorems regarding local stability used in this paper are relegated to the Appendix A. In this subsection, we first give the Jacobian of the CH-NTP dynamic at the MPE and then analyze the stability near the DUE.
@@ 3.2.1 Jacobian matrix
With Lemma 2 established, we can now derive the Jacobian matrix for the CH-NTP dynamic and utilize it to analyze local stability at any MPE by examining its eigenvalues.
@@
where $D_{k}\equiv D\left[\boldsymbol{\pi}^{k,(t+1)}\right]$ representing the Jacobian of the route cost functions evaluated at $\boldsymbol{\pi}^{k,(t+1)}$; $Q^{k}\equiv Q\left[\boldsymbol{x}^{k,(t)}-\gamma\boldsymbol{c}\left(\boldsymbol{\pi}^{k,(t+1)}\right)\right]$; $I$ the identity matrix of size $\sum_{w}{|R_{w}|}$; $\hat{Q}_{1}^{0}\equiv Q\left[q_{1}^{0}\tilde{\boldsymbol{x}}^{(t)}-\hat{\gamma}\boldsymbol{c}\left(\boldsymbol{\pi}^{0,(t+1)}\right)\right]$; $\hat{Q}_{2}^{0}\equiv Q\left[q_{2}^{0}\tilde{\boldsymbol{x}}^{(t)}-\hat{\gamma}\boldsymbol{c}\left(\boldsymbol{\pi}^{0,(t+1)}\right)\right]$; and $\hat{Q}_{2}^{1}\equiv Q\left[q_{2}^{1}\tilde{\boldsymbol{x}}^{(t)}-\hat{\gamma}\boldsymbol{c}\left(\boldsymbol{\pi}^{1,(t+1)}\right)\right]$.
@@ 3.2.2 Stability around the DUE
We are particularly interested in how the travelers’ prediction behaviors would affect the stability near the DUE. Numerically, it is always tractable to calculate $JP$’s eigenvalues at the DUE and check its local stability. However, general analytical insights on the parametric space may rely on certain assumptions. Recognizing this, we assume that as the day-to-day process evolves, the projection always yields a positive flow on each route. Under this assumption, the NTP dynamic reduces to a closed-form expression (Sandholm 2010, Xiao, Yang, and Ye 2016). The $\alpha$ and $\gamma$ (and $\hat{\alpha}$ and $\hat{\gamma}$) can be simply combined (specifically, multiplied) to represent the sensitivity to the cost difference between two routes; see Appendix A in Xiao, Yang, and Ye (2016). We can therefore fix $\alpha=\hat{\alpha}=1$ and vary $\gamma$ and $\hat{\gamma}$ while the behavioral explanation is not compromised.
We divide the discussion into two parts: (i) when higher-step travelers can exactly predict lower-step travelers’ switching tendency, i.e., $\hat{\gamma}=\gamma$; and (ii) when the prediction is inaccurate, i.e., $\hat{\gamma}\neq\gamma$. The following assumption is made on the route cost function.
@@ 3.2.2.1   Perfect prediction $\hat{\gamma}=\gamma$
For a dynamical system with Eqs. (9)-(12) featuring any $|K|\in\{1,2,3\}$, if the projection operator can always generate positive flow on each route (hence $\hat{\alpha}$ and $\alpha$ can be set to 1) and $\hat{\gamma}=\gamma$, its DUE under Assumption 1 is locally stable if all the moduli of the eigenvalues of matrix $\bar{Q}(I-\gamma D_{\ast})$ are less than 1, where $D_{\ast}$ is the Jacobian of the route cost function evaluated at the aggregate DUE, $\tilde{\boldsymbol{x}}^{\ast}$, and $\bar{Q}$ is defined in Lemma 2.
@@
The key distinction between models with different $|K|$ values appears in their response to instability. Our numerical experiments show that when instability occurs with $|K|=1$, the system tends to permanently oscillate near the DUE, repeatedly crossing this equilibrium point. Conversely, systems with $|K|\geq 2$ are more likely to evolve toward and remain at an MPE that differs from the DUE.
@@
Please see Appendix E. ∎
Remark 3.
To provide realistic context for our stability conditions, note that the critical threshold $\bar{\gamma}\equiv\frac{2}{\max_{i}{\mu_{i}}}$ reflects the network’s inherent properties, specifically how steeply costs increase with traffic flow. Networks with steeper cost functions have larger eigenvalues $\mu_{i}$ of $\bar{Q}D_{\ast}$, resulting in a smaller $\bar{\gamma}$ and making them naturally more prone to instability. The parameter $\gamma$ represents travelers’ sensitivity to cost differences between routes. When $\gamma<\bar{\gamma}$, the system maintains stability at equilibrium, while larger values of $\gamma$ may lead to oscillatory traffic patterns as travelers rapidly shift between routes.
@@
Although these findings emerge from a simple two-route network analysis, they generalize to arbitrary networks. This generalization is formally established in the following proposition.
@@
Consider a dynamical system with Eqs. (9)-(11) with $|K|=2$ where the projection can always generate positive flow on each route (hence $\hat{\alpha}$ and $\alpha$ can be set to 1). Denote $\beta_{i}$ as $\bar{Q}D_{\ast}$’s $i$-th eigenvalue and $\beta_{i}\geq 0,\forall i$. The DUE is locally stable if and only if $\left|{\gamma\hat{\gamma}\beta_{i}^{2}-2\gamma\beta_{i}+1}\right|<1,\forall i$.
@@
3.3 Analyses of non-DUE MPE (MPE that are not corresponding to DUE)
This section analyzes the properties and implications of MPE that do not correspond to DUE. Section 3.3.1 characterizes a fundamental property of non-DUE MPE that each traveler class must abandon at least one route due to strategic anticipation of congestion. This property suggests that non-DUE MPE are more likely to emerge when travelers extensively anticipate others’ behaviors. We confirm this hypothesis in Section 3.3.2 through numerical experiments. Section 3.3.3 examines the total network cost implications of these equilibria.
3.3.1 Characterization of non-DUE MPE
Proposition 5.
Consider the CH-NTP dynamic of Eqs. (9)-(11) with $|K|=2$. At any non-DUE MPE, denoted as $\boldsymbol{x}^{\dagger}\equiv(\boldsymbol{x}^{k,\dagger},\boldsymbol{x}^{k,\dagger}\in\Omega_{p^{k}},k=0,1)^{T}$, each traveler class has at least one unused route $j(k)$ where $x_{j(k)}^{k,\dagger}=0$, despite the original DUE admitting strictly positive flow on each route.
Proof.
Please see Appendix G.∎
Remark 5.
Intuitively, this abandonment stems from travelers’ extensive prediction capabilities: higher-step travelers avoid seemingly costless routes due to their reasoning that these routes would become congested through widespread adoption by lower-step travelers; this creates a strategic redistribution where higher-step travelers select longer routes to avoid predicted congestion. These patterns become self-reinforcing once established: persistent over-predictions of lower-step travelers’ responses confines higher-step travelers to some routes, even as actual costs remain elevated on these chosen routes.
3.3.2 When non-DUE MPE emerge?
From the analysis in Remark 5, we expect non-DUE MPE to emerge only when $\hat{\gamma}$ exceeds a certain threshold. This is confirmed through numerical experiments on a parallel three-route network from Wang, Guo, and Huang (2021), using CH-NTP dynamics with $|K|=2$, $p^{0}=p^{1}=0.5$ and $\hat{\alpha}=\alpha=1$. (We have conducted additional tests across numerous parameter configurations, and the general patterns and conclusions remain consistent.)
Figures 1(a)-1(d) depict the evolution trajectories of aggregate route flows from different initial points under increasing $\hat{\gamma}$ values. Figure 1(a) shows that with a small $\hat{\gamma}(0.8)$, all initial points converge to the DUE. However, as $\hat{\gamma}$ increases to 1.3, 1.8, and 2.5 (Figures 1(b)-1(d)), multiple non-DUE MPE emerge, and the probability of evolution toward these non-DUE MPE increases. At $\hat{\gamma}=2.5$, all trajectories evolve exclusively toward non-DUE MPE.
(a) $\hat{\gamma}=0.8$
(b) $\hat{\gamma}=1.3$
(c) $\hat{\gamma}=1.8$
(d) $\hat{\gamma}=2.5$
Figure 1: Evolution trajectories of aggregate route flows under varying $\hat{\gamma}$ values for CHT-NTP dynamics with $|K|=2$, $p^{0}=p^{1}=0.5$, and $\hat{\alpha}=\alpha=1$ in the parallel three-route network from Wang, Guo, and Huang (2021). Each trajectory originates from distinct boundary points where at least one route having zero flow.
3.3.3 Total network cost at non-DUE MPE
When the system settles at a non-DUE MPE, does it result in higher or lower total network costs compared to DUE? The answer proves to be nuanced and depends critically on the network parameters. This complexity becomes evident even in the perhaps simplest case of a two-route parallel network with linear cost functions.
Example 1.
Consider the CH-NTP dynamic with Eqs. (9)-(11) with $|K|=2$ in a two-parallel-route network with linear cost functions $c_{1}(\tilde{x}_{1})=a_{1}\tilde{x}_{1}+b_{1}$ and $c_{2}(\tilde{x}_{2})=a_{2}\tilde{x}_{2}+b_{2}$. At the DUE, both routes exhibit positive aggregate flows. There exist two non-DUE MPE: one is $\boldsymbol{x}^{\dagger}\equiv(x_{1}^{0,\dagger},x_{2}^{0,\dagger},x_{1}^{1,\dagger},x_{2}^{1,\dagger})=(p^{0}d,0,0,p^{1}d)^{T}$, and the other is $\boldsymbol{x}^{\dagger}=(0,p^{0}d,p^{1}d,0)^{T}$. The non-DUE MPE produces lower total costs than DUE when either:
  • (i)
    The dynamic finally settles at $\boldsymbol{x}^{\dagger}=(p^{0}d,0,0,p^{1}d)^{T}$, $b_{1}<b_{2}$ and $p^{0}\in\left(\frac{a_{2}}{a_{1}+a_{2}},\frac{b_{2}-b_{1}+a_{1}d}{(a_{1}+a_{2})d}\right)$; or
  • (ii)
    The dynamic finally settles at $\boldsymbol{x}^{\dagger}=(0,p^{0}d,p^{1}d,0)^{T}$, $b_{1}>b_{2}$ and $p^{1}\in\left(\frac{b_{2}-b_{1}+a_{2}d}{(a_{1}+a_{2})d},\frac{a_{2}}{a_{1}+a_{2}}\right)$.
Proof.
Please see Appendix H. ∎
Example 1 reveals that when the proportion of strategic thinkers falls within specific critical ranges that depend on the cost function parameters, the system may settle at a non-DUE MPE with flow pattern that favor routes with lower marginal costs rather than average costs. This phenomenon resembles aspects of system optimal behavior yet emerges endogenously from cognitive hierarchy dynamics. In all other cases not covered by the conditions in Example 1, strategic thinking becomes detrimental to overall system efficiency.